Solve the differential equation y'=x(1+y^2)
1
Expert's answer
2019-10-17T09:04:21-0400

"y^{\prime}=x(1+y^2)\\
\frac{dy}{1+y^2}=x\,dx\\
\arctan(y) = \frac{x^2}{2}+C\\
y=\tan(\frac{1}{2}(x^2+C))"